# Evaluating Concept-based Explanations of Language Models: A Study on Faithfulness and Readability

### v2

@@ Abstract
Despite the surprisingly high intelligence exhibited by Large Language Models (LLMs), we are somehow intimidated to fully deploy them into real-life applications considering their black-box nature. Concept-based explanations arise as a promising avenue for explaining what the LLMs have learned, making them more transparent to humans. However, current evaluations for concepts tend to be heuristic and non-deterministic, e.g. case study or human evaluation, hindering the development of the field. To bridge the gap, we approach concept-based explanation evaluation via faithfulness and readability. We first introduce a formal definition of concept generalizable to diverse concept-based explanations. Based on this, we quantify faithfulness via the difference in the output upon perturbation. We then provide an automatic measure for readability, by measuring the coherence of patterns that maximally activate a concept. This measure serves as a cost-effective and reliable substitute for human evaluation. Finally, based on measurement theory, we describe a meta-evaluation method for evaluating the above measures via reliability and validity, which can be generalized to other tasks as well. Extensive experimental analysis has been conducted to validate and inform the selection of concept evaluation measures. 111Codes available at https://github.com/hr-jin/Concept-Explaination-Evaluation
@@ 1 Introduction
Explainable Artificial Intelligence (XAI) holds significant value in pre-trained language models’ mechanism understanding Li et al. (2022), performance enhancement Wu et al. (2023); Ribeiro et al. (2016), and security Burger et al. (2023); Zou et al. (2023). Previous XAI endeavors concentrated on attribution methods Lundberg and Lee (2017); Sundararajan et al. (2017). The attribution methods identify “where” the model looks rather than “what” it comprehends Colin et al. (2022), typically offering local explanations for a limited number of input samples, restricting their utility in practical settings Colin et al. (2022); Adebayo et al. (2018). Concept-based explanations Kim et al. (2018); Hennigen et al. (2020); Cunningham et al. (2023); Fel et al. (2023b) can mitigate the limitations of attribution methods by recognizing high-level Kim et al. (2018) patterns (see Fig. 1), which provide concise, human-understandable explanations of models’ internal state.
@@
First, we provide a unified definition of diverse concept-based explanation methods and quantify faithfulness under this formalization (C1). By summarizing common patterns of concept-based explanation, we provide a formal definition of a concept, which can generalize to both supervised and unsupervised, post-hoc and interpretable-by-design methods, language and vision domains. Based on this, we quantify a concept’s faithfulness via the difference in the output caused by a perturbation of the concept. Furthermore, we approach the proper degree of perturbation of different representations via an optimization problem.
@@
Third, we describe a meta-evaluation method for evaluating the evaluation measures based on measurement theory (C3). Measurement theory Allen and Yen (2001); Xiao et al. (2023) has been long utilized to verify whether a measurement is reliable and valid. Approaching via reliability and validity, this meta-evaluation method is useful for evaluating the measures for concepts and can be generalized to analyze the effectiveness of other measures, for example, measures for other types of explanations and other natural language tasks.
@@ 2 Concept Formalization
In this paper, we primarily focus on explaining LLMs as black-box models. Meanwhile, our method can generalize to many other deep classification models, including image models (see Appx.C). As illustrated in Fig. 1, we consider the black-box model to take an input from a dataset and output , a -class classification probability distribution. In text generation, is the vocabulary size. For the -th layer to be interpreted, given a sequence of tokens , their hidden representations are . The output classification logits are .
Within the context of Deep Neural Networks (DNNs), we summarize common patterns of concepts and establish a unified framework. Specifically, each concept is represented as a virtual neuron defined by an activation function that maps a hidden representation into a real value , where a positive output signifies activation. For a human-understandable concept, we may provide a semantic expression, e.g., terminologies related to computer networks (Fig. 1), by summarizing patterns that maximally activate it.
@@
| ProtoPNet* Chen et al. (2019) | image | ||
| Unupervised | NetDissect Bau et al. (2017) | image | |
@@
Our formalization can integrate diverse concept explanation methods, as shown in Tab. 1. This includes supervised and unsupervised methods, post-hoc and interpretable-by-design methods, and is also applicable to image backbone models.
@@ 3.1 Faithfulness
with being the probability distributions of output vocabulary.
Concept perturbation. Based on the formalization of concepts in Sec. 2, we view this problem as an optimization problem. As the concept formalization provided above encapsulates diverse kinds of concepts, this transformation allows the perturbation strategies to generalize beyond the linear form of concepts, like Chen et al. (2019).
@@ 3.2 Readability
Readability assesses the extent to which humans can comprehend the extracted concept Lage et al. (2019). Most of the time, when patterns that maximally activate a concept are coherent (see example in Fig. 1), can the concept be easily understandable to humans. As cross-sample patterns are extracted from a large corpus, diverse samples are needed to evaluate its readability. Although previous efforts have made some progress in evaluating readability, they confront the challenge of ensuring data comprehensiveness while minimizing cost. Tab. 2 compares different measures for readability, including human evaluation Kim et al. (2018); Ghorbani et al. (2019), LLM-based measures Bills et al. (2023); Singh et al. (2023), and our proposed coherence-based measures.
@@
Human evaluation. Existing approaches predominantly rely on case studies and user studies Kim et al. (2018); Ghorbani et al. (2019); Chen et al. (2019), asking humans to score a concept given a limited number of demonstrative samples. They are subject to issues of validation, standardization, and reproducibility Clark et al. (2021); Howcroft et al. (2020).
LLM-based. As inexpensive human substitutes, LLMs have been utilized in evaluating concept-based explanations. A typical LLM-based score Bills et al. (2023); Singh et al. (2023) is obtained by: 1) letting LLM summarize a natural language explanation for the concept (e.g., semantic expression in Fig. 1) given formatted samples that maximally activates on the concept and activations ; 2) letting LLM guess the activation given only sample text and the generated explanation; 3) calculating an explanation score based on the variance between true activation and the simulated activation. However, the number of samples inputted to LLMs (4 in Bills et al. (2023)) in step 1 is limited to maximum input length. This limits the comprehensiveness of the generated explanation, as shown in a case study in Appx. D. Even if the maximum input length is extended to 200k+ like Claude 3111www.anthropic.com/news/claude-3-family, it may suffer from high computation cost and poor performance in long-dependency tasks Li et al. (2023).
@@
Probabilities () are estimated based on word (co-)occurrence counts in the corpus. To prevent zero values in logarithmic operations, a small value is introduced. embeds a word to a continuous semantic space, for example, using embedding models like BERT.
@@ 4 Meta Evaluation
The potential absence of a golden label for extracted concepts, compounded by the vast and often infinite array of possible concept sets Hoffman et al. (2018), complicates assessment. Borrowing metrics from measurement theory Allen and Yen (2001) and psychometrics Wang et al. (2023b), our meta-evaluation focus centers on reliability and validity, guided by the methodological framework outlined in Xiao et al. (2023). Our meta-evaluation methods can also be generalized to measures of a broader scope, including other XAI methods and other natural language tasks like generation.
@@ 4.1 Reliability
where is the observed score, and is the true score, denotes variance of a random variable . Typically, the minimal standard for an acceptable measure is 0.9 Nunnally and Bernstein (1994).
@@
Inter-rater reliability measures the degree of agreement across raters, calculated as score correlation among them. In this paper, we apply Kendall’s Kendall (1938) to measure pairwise correlation among two raters using a scale that is ordered:
@@ 4.2 Validity
Validity is crucial in assessing how well a test measures the intended construct Nunnally and Bernstein (1994), be it readability or faithfulness in our case. We focus on concurrent validity, evaluating the extent to which a test score predicts outcomes on a validated measure Cronbach and Meehl (1955), and construct validity, examining how well indicators represent an unmeasurable concept Cronbach and Meehl (1955). Construct validity can be further divided into convergent validity and divergent validity.
@@
Convergent validity verifies whether measures of the same construct are indeed related. For example, if the purposed faithfulness measures are related. As the underlying construct is often inaccessible to directly assess the measures’ concurrent validity, convergent validity provides a statistic tool to assess construct validity via its relation Kendall (1938) with other measures of the same construct.
@@ 5.2.2 Validity
Convergent Validity is inspected via correlation between measures of the same construct. Faithfulness measures (A) displayed moderate correlations in general, averaging around 0.5. Aggrement between measures with the same perturbation strategy or difference measurement is higher than others, indicating their potential relation. *-TClass and *-PClass showed a higher correlation, due to the consistency between prediction and true classes in well-trained language models. In the meantime, the agreement of readability measures (B) on either the input side or output side is moderate.
@@ 5.3 Comparison of Explanation Methods
We conducted a comparative assessment of three different baseline methods on the language domain, including the concepts of neuron Bills et al. (2023), sparse autoencoder Cunningham et al. (2023), and TCAV Kim et al. (2018). The results for both the neuron and sparse autoencoder were computed as the average values across 100 randomly sampled concepts from the concept set, whereas the TCAV concept was trained on a corpus related to LLM’s harmful QA Bhardwaj and Poria (2023), yielding a singular concept. The results of this analysis are shown in Fig. 4.
@@ 7 Limitations
Topic coherence is not designed to be the ultimate or perfect solution for measuring readability. Other aspects of readability, such as meaningfulness Ghorbani et al. (2019), may also be worth exploring. In the future, we are interested in investigating how these aspects could be quantified automatically, building a more comprehensive landscape of readability.
Constrained by limited GPU resources, we used smaller versions of LLM, focusing primarily on the 3rd layer of Pythia-70M for our analysis. Due to budget constraints, our evaluation of the LLM-Score was restricted to 200 concepts, incurring a cost of around $1 for a single concept. While this setup, on par with Cunningham et al. (2023) and more general than Bricken et al. (2023), allowed for fast analysis and comparison with existing literature, expanding our analysis to larger models could yield more insightful conclusions in the future.
8 Ethical Statements
@@ Appendix F Implementation Details
Following Bills et al. (2023); Cunningham et al. (2023), we adopt the extraction of neuron activation as the output of the MLP layer in each layer, where each dimension corresponds to a neuron. Specifically, for a feed-forward layer , the MLP output/neurons are . Furthermore, the disentanglement-based baseline can utilize these extracted neurons as inputs to discover mono-semantic concepts, leveraging sparse autoencoders. For TCAV Kim et al. (2018) baseline, we train the CAV on harmful QA dataset Bhardwaj and Poria (2023)
@@
For LLM-based readability score Bills et al. (2023), we adopt this adjusted algorithm with gpt-4-turbo-preview as the simulator, due to new limitations in calculating logprobs on the input side. When extracting patterns that maximally activate a concept, we keep only the top 10 tokens with the largest activation or contribution to high-activation tokens.

### v3

# Evaluating Readability and Faithfulness of Concept-based Explanations

@@ Abstract
With the growing popularity of general-purpose Large Language Models (LLMs), comes a need for more global explanations of model behaviors. Concept-based explanations arise as a promising avenue for explaining high-level patterns learned by LLMs. Yet their evaluation poses unique challenges, especially due to their non-local nature and high dimensional representation in a model’s hidden space. Current methods approach concepts from different perspectives, lacking a unified formalization. This makes evaluating the core measures of concepts, namely faithfulness or readability, challenging. To bridge the gap, we introduce a formal definition of concepts generalizing to diverse concept-based explanations’ settings. Based on this, we quantify the faithfulness of a concept explanation via perturbation. We ensure adequate perturbation in the high-dimensional space for different concepts via an optimization problem. Readability is approximated via an automatic and deterministic measure, quantifying the coherence of patterns that maximally activate a concept while aligning with human understanding. Finally, based on measurement theory, we apply a meta-evaluation method for evaluating these measures, generalizable to other types of explanations or tasks as well. Extensive experimental analysis has been conducted to inform the selection of explanation evaluation measures. 111Codes available at https://github.com/hr-jin/Concept-Explanation-Evaluation
@@ 1 Introduction
Explainable Artificial Intelligence (XAI) holds significant value in pre-trained language models’ mechanism understanding Li et al. (2022), visualization Yang et al. (2024), performance enhancement Wu et al. (2023b); Ribeiro et al. (2016); Wang et al. (2022), and security Burger et al. (2023); Zou et al. (2023). Previous XAI algorithms have been applied to NLP tasks Wu et al. (2023a), vision tasks Wang et al. (2023b) and recommendation Jin et al. (2022); Yang et al. (2022). These include natural language explanation Zhang et al. (2024); Lee et al. (2022), attention explanation Chen et al. (2019b); Gao et al. (2019), and especially attribution methods Lundberg and Lee (2017); Sundararajan et al. (2017); Guan et al. (2019). The attribution methods identify “where” the model looks rather than “what” it comprehends Colin et al. (2022), typically offering local explanations for a limited number of input samples, restricting their utility in practical settings Colin et al. (2022); Adebayo et al. (2018). Concept-based explanations Kim et al. (2018); Cunningham et al. (2023); Fel et al. (2023b) can mitigate the limitations of attribution methods by recognizing high-level Kim et al. (2018) patterns (see Fig. 1), which provide concise, human-understandable explanations of models’ internal state.
@@
First, we provide a unified definition of diverse concept-based explanation methods and quantify faithfulness under this formalization (C1). By summarizing common patterns of concept-based explanation, we provide a formal definition of a concept, which can generalize to both supervised and unsupervised, post-hoc and interpretable-by-design methods, language and vision domains. Based on this, we quantify the faithfulness of a concept explanation via perturbation. We ensure adequate perturbation in the high-dimensional space for different concepts via an optimization problem.
@@
Third, we apply the classic measurement theory to perform a meta-evaluation on the faithfulness and readability measures (C3). Measurement theory Allen and Yen (2001); Xiao et al. (2023) has been long utilized to verify whether a measurement is reliable and valid. Approaching via reliability and validity, this meta-evaluation method is useful for evaluating the measures for concepts and can be generalized to analyze the effectiveness of other measures, for example, measures for other types of explanations and other natural language tasks. Experimental results have filtered out 4 measures with low reliability, i.e. LLM-Score, GRAD-Loss, IN-UCI, IN-UMass, and verified the remaining faithfulness and readability measures’ validity.
@@ 2 Concept Formalization
In this paper, we primarily focus on explaining LLMs as black-box models. Meanwhile, our method can be generalized to many other deep classification models, including image models (see Appx.C). As illustrated in Fig. 1, we consider the black-box model to take an input from a dataset and output , a -class classification result. In text generation, is the vocabulary size. For the -th layer to be interpreted, given a sequence of input tokens , their corresponding hidden representations are . The output classification logits are .
Within the context of Deep Neural Networks (DNNs), we summarize common patterns of concepts and establish a unified framework. Specifically, each concept is represented as a virtual neuron defined by an activation function that maps a hidden representation into a real value , where a positive output signifies activation. For each concept, a semantic expression may be given by humans or LLMs, depending on the concept explanation methods Kim et al. (2018); Bills et al. (2023). Some methods take concepts and semantic expressions predefined by humans as inputs (e.g., Kim et al. (2018)), while others require additional steps to produce a semantic expression based on highly activated tokens and samples of the extracted concepts Bills et al. (2023). Specifically, given high-activation samples of the concept and the highly activated tokens in these samples (e.g., “Internet, computer, networks, …”), an LLM or a human labeler provides a semantic expression that summarizes their common patterns (e.g., “Terminologies related to computer networks”) Bills et al. (2023).
@@
| ProtoPNet* Chen et al. (2019a) | image | ||
| Unsupervised | NetDissect Bau et al. (2017) | image | |
@@
Our formalization can integrate diverse concept explanation methods, as shown in Tab. 1. This includes both supervised methods that require prior information about concepts (e.g., input samples that contain and do not contain the concepts) Kim et al. (2018) and unsupervised methods that do not rely on such prior information Ghorbani et al. (2019). Our method also works for both post-hoc explanation methods that interpret a model after it is trained Kim et al. (2018) and interpretable-by-design approaches that integrate interpretability mechanisms directly into the model’s architecture before training Koh et al. (2020). Additionally, it applies to image backbone models as well.
@@ 3.1 Faithfulness
with being the probability distribution of output vocabulary.
Concept perturbation. Based on the formalization of concepts in Sec. 2, we view this problem as an optimization problem. As the concept formalization provided above encapsulates diverse kinds of concepts, this transformation allows the perturbation strategies to generalize beyond the linear form of concepts, like Chen et al. (2019a).
@@ 3.2 Readability
Readability assesses the extent to which humans can comprehend the extracted concept Lage et al. (2019). Most of the time, when patterns that maximally activate a concept are coherent (see example in Fig. 1), can the concept be easily understandable to humans. We design coherence measures based on OpenAI’s pipeline Bills et al. (2023) for human evaluation of concept quality. They presented human labelers with fragments where highly activated tokens were shown with color highlighting and asked the humans to try summarizing the commonalities of these highly activated tokens. We automate this process by assessing the commonality of highly activated tokens via co-occurrence or embedding similarity.
As cross-sample patterns are extracted from a large corpus, diverse samples are needed to evaluate a concept’s readability. Although previous efforts have made some progress in evaluating readability, they confront the challenge of ensuring data comprehensiveness while minimizing cost. Tab. 2 compares different measures for readability, including human evaluation Kim et al. (2018); Ghorbani et al. (2019), LLM-based measures Bills et al. (2023); Singh et al. (2023), and our proposed coherence-based measures. For the LLM-based evaluation, we considered Bills et al. (2023); Bricken et al. (2023), which used less than 100 samples. For human evaluation, we considered the classical method by Ghorbani et al. (2019), where each human rater scored no more than 20 samples per concept.
@@
Human evaluation. Existing approaches predominantly rely on case studies and user studies Kim et al. (2018); Ghorbani et al. (2019); Chen et al. (2019a), asking humans to score a concept given a limited number of demonstrative samples. They are subject to issues of validation, standardization, and reproducibility Clark et al. (2021); Howcroft et al. (2020).
LLM-based. As inexpensive human substitutes, LLMs have been utilized in evaluating concept-based explanations. A typical LLM-based score Bills et al. (2023); Singh et al. (2023) is obtained by: 1) letting LLM summarize a natural language explanation for the concept (e.g., semantic expression in Fig. 1) given formatted samples that maximally activates on the concept and activations ; 2) letting LLM guess the activation given only sample text and the generated explanation; 3) calculating an explanation score based on the variance between true activation and the simulated activation. However, the number of samples inputted to LLMs (4 in Bills et al. (2023)) in step 1 is limited to maximum input length. This limits the comprehensiveness of the generated explanation, as shown in a case study in Appx. D. Even if the maximum input length is extended to 200k+ like Claude 3222www.anthropic.com/news/claude-3-family, it may suffer from high computation cost and poor performance in long-dependency tasks Li et al. (2023).
@@
Probabilities are estimated based on word occurrence frequency in the corpus. To prevent zero values in logarithmic operations, a small value is introduced. embeds a word to a continuous semantic space, for example, using embedding models like BERT.
@@ 4 Meta Evaluation
How can we discern the effectiveness among possible measures available for evaluating concept-based explanations? Borrowing metrics from measurement theory Allen and Yen (2001) and psychometrics Wang et al. (2023c); Xiao et al. (2023), our meta-evaluation focus centers on reliability and validity, guided by the methodological framework outlined in Allen and Yen (2001). Our meta-evaluation methods can generalized to measures of a broader scope, including other XAI methods and other natural language tasks like generation.
@@ 4.1 Reliability
where is the observed score, and is the true score, denotes the variance of a random variable . Typically, the minimal standard for an acceptable measure is 0.9 Nunnally and Bernstein (1994).
@@
Inter-rater reliability measures the degree of agreement across raters, calculated as score correlation among them. In this paper, we apply Kendall’s Kendall (1938) to measure pairwise correlation among raters using a scale that is ordered:
@@ 4.2 Validity
Validity is crucial in assessing how well a test measures the intended construct Nunnally and Bernstein (1994). A construct refers to the underlying criterion to be measured. In our case, it is faithfulness or readability. We focus on concurrent validity, evaluating the extent to which a test score predicts outcomes on a validated measure Cronbach and Meehl (1955), and construct validity, examining how well indicators represent an unmeasurable concept Cronbach and Meehl (1955). Construct validity can be further divided into convergent validity and divergent validity.
@@
Convergent validity verifies whether measures of the same construct are indeed related. For example, whether the purposed faithfulness measures are related to each other. As the underlying construct is often inaccessible to directly assess the measures’ concurrent validity, convergent validity provides a statistic tool to assess construct validity via its relation Kendall (1938) with other measures of the same construct.
@@ 5.2.2 Validity
Convergent Validity is inspected via correlation between measures of the same construct. Faithfulness measures (A) displayed moderate correlations in general, averaging around 0.5. Agreement between measures with the same perturbation strategy or difference measurement is higher than others, indicating their potential relation. *-TClass and *-PClass showed a higher correlation, due to the consistency between prediction and true classes in well-trained language models. In the meantime, the agreement of readability measures (B) on either the input side or output side is moderate.
@@ 5.3 Comparison of Explanation Methods
We conducted a comparative assessment of three different baseline methods on the language domain, including the concepts of neuron Bills et al. (2023), sparse autoencoder Cunningham et al. (2023), and TCAV Kim et al. (2018). The results for both the neuron and sparse autoencoder were computed as the average values across 100 randomly sampled concepts from the concept set. We derive the supervised concept using TCAV following Kim et al. (2018); Xu et al. (2024). Initially, LLM’s harmful QA Bhardwaj and Poria (2023) is treated as positive examples, and random texts are treated as negative examples. Their hidden representations are used to train a linear classifier, which aims to differentiate the representations of positive examples from negative ones. The trained classifier is treated as the concept’s activation function. The results of this analysis are shown in Fig. 4.
@@
Besides, we found that LLM has learned a seemingly redundant yet interesting pattern for the first concept shown for sparse autoencoder (e.g., north,·west, ·east, ·South, ·North, South, ·northern). Though these tokens are quite similar for humans, we do not know whether they are considered the same for LLMs. The embedding similarity between these tokens reflects LLMs’ ability to model them just like how humans perceive them as similar.
@@ 6 Conclusion
Topic coherence is not designed to be the ultimate or perfect solution for measuring readability. Other aspects of readability, such as meaningfulness Ghorbani et al. (2019), may also worth exploring. In the future, we are interested in investigating how these aspects could be quantified automatically, building a more comprehensive landscape of readability.
Due to limited GPU resources and budget constraints, we used smaller versions of LLM, focusing primarily on the 3rd layer of Pythia-70M for our analysis. And our evaluation of the LLM-Score was restricted to 200 concepts, incurring a cost of around $1 for a single concept. While this setup, on par with Cunningham et al. (2023) and more general than Bricken et al. (2023), allowed for fast analysis and comparison with existing literature, expanding our analysis to larger models could yield more insightful conclusions in the future.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) (NO. 62476279), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, Kuaishou, and the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China No. 24XNKJ18. This work was partially done at Beijing Key Laboratory of Big Data Management and Analysis Methods and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education. This research was supported by Public Computing Cloud, Renmin University of China.
Ethical Statements
@@ Appendix F Implementation Details
Following Bills et al. (2023); Cunningham et al. (2023), we adopt the extraction of neuron activation as the output of the MLP layer in each layer, where each dimension corresponds to a neuron. Specifically, for a feed-forward layer , the MLP output/neurons are . Furthermore, the disentanglement-based baseline can utilize these extracted neurons as inputs to discover mono-semantic concepts, leveraging sparse autoencoders. We obtain the concept activation function of TCAV following Kim et al. (2018). We treat LLM’s harmful QA Bhardwaj and Poria (2023) as positive examples, and random texts as negative examples. Then, a linear classifier is trained on their hidden representations to classify harmful examples. The trained classifier’s output function is regarded as the concept’s activation function.
@@
For LLM-based readability score Bills et al. (2023), we follow OpenAI’s pipeline as illustrated in Bills et al. (2023). We show a detailed prompt for generating an explanation/semantic expression of a concept based on its activation in Fig. 9. We adopt this adjusted algorithm with gpt-4-turbo-preview as the simulator, due to new limitations in calculating logprobs on the input side. When extracting patterns that maximally activate a concept, we keep only the top 10 tokens with the largest activation or contribution to high-activation tokens.
@@ Appendix G User study settings
Given input or output side tokens for a concept, each of our human labelers gives one readability score by simultaneously considering the three aspects, including semantic, grammatical or syntactic, and morphological information. More specifically, a concept is considered highly readable, if it is related to a specific topic such as computer systems (semantically interesting), is associated with a specific grammar or syntax (grammatically or syntactically interesting), or consists of tokens that share a similar structure or a form such as all being usable as suffixes for a certain token (morphologically interesting).